OpenStudy (anonymous):

This is my first time using this web site. Could anyone help me with the solution to [3x^2/3y^1/4]^3 i dont know if i am just over complicating the simple or what. any help would be greatly appreciated.

OpenStudy (anonymous):

thank you for your prompt reply. that said i guess i wrote the problem wrong.

OpenStudy (anonymous):

the x is to the two thirds power(2/3) and the y is just y to the one fourth power(1/4) ...please forgive me if i wrote it wrong.

OpenStudy (anonymous):

oh yeah also it is not x over y it is x times y .

OpenStudy (anonymous):

how did you draw?

OpenStudy (unklerhaukus):

\[\left[3x^{2/3} y^{1/4}\right]^3 \]?

OpenStudy (unklerhaukus):

\[\left[ab^nc^d\right]^f=\left[a^fb^{n\times f}c^{d\times f}\right]\]

OpenStudy (anonymous):

Yes sir your first response is my problem. I am having Internet issues today.

OpenStudy (precal):

see the blue middle key that states "Draw", press that and use your mouse to draw

OpenStudy (precal):

seems I did try to help you after all

OpenStudy (precal):

@precal will send me a notification were you able to figure this one out? Let me know and I will look at it again.

OpenStudy (anonymous):

I was not able to figure it out. @precal

OpenStudy (precal):

what page number and problem, easier for me to look it up

OpenStudy (anonymous):

It's page 2 problem fifteen of the packet.

OpenStudy (precal):

ok let me look at it

OpenStudy (precal):

 ok apply your exponent 3 to all of your bases